Find the angle of the largest right circular cone which can be inscribed in a sphere of

radius 9 inches.


1
Expert's answer
2022-05-17T14:43:28-0400


Let "R=" the radius of the sphere.

Let "AD=r, BD=h, \angle ABC=\theta."


"\triangle ABC"

"AC=2r, AO=OC=R, \angle AOC=2\angle ABC=2\theta"

The Law of Cosines



"(2r)^2=R^2+R^2-2R^2\cos(2\theta)""4r^2=2R^2(2\sin^2\theta)""r=R\sin \theta""h=r\tan(\theta/2)=R\sin \theta\tan(\theta/2)""V_{cone}=\dfrac{1}{3}\pi r^2 h""V_{cone}=V_{cone}(\theta)=\dfrac{1}{3}\pi R^3\sin^3 \theta\tan(\theta/2)""(V_{cone})'_{\theta}=\dfrac{1}{3}\pi R^3\bigg(3\sin^2 \theta\cos(\theta)\tan(\theta/2)""+\dfrac{1}{2}\sin^3\theta(\dfrac{1}{\cos ^2(\theta/2)})\bigg)"

Find the critical numbr(s)



"(V_{cone})'_{\theta}=0""\dfrac{1}{3}\pi R^3\bigg(3\sin^2 \theta\cos\theta\tan(\theta/2)""+\dfrac{\sin^3\theta}{2\cos ^2(\theta/2)}\bigg)=0""\dfrac{\sin^3 \theta(3\cos\theta+1)}{2\cos ^2(\theta/2)}=0""\cos \theta=-\dfrac{1}{3}"

If "0< \theta<\pi-\cos^{-1}(1/3), (V_{cone})'_{\theta}>0, V_{cone}" increases.


If "\pi-\cos^{-1}(1/3)<\theta<\pi, (V_{cone})'_{\theta}<0, V_{cone}" decreases.


The volume of inscribed cone has the absolute maximum at


"\theta=\pi-\cos^{-1}(1/3)"